OpenStudy (anonymous):

What is the value of 7P3 ?

OpenStudy (doc.brown):

Does P mean permutation?\[\frac{7!}{(7-3)!}\]\[\frac{7!}{4!}\]\[\frac{7\times 6\times 5\times 4\times 3\times 2\times 1}{4\times 3\times 2\times 1}\]\[\frac{7\times 6\times 5\times \cancel{4}\times \cancel{3}\times \cancel{2}\times 1}{\cancel{4}\times \cancel{3}\times \cancel{2}\times 1}\]\[=7\times 5\times 4\]